OpenStudy (anonymous):

Four of Ethan's family members went on a 3 and one-third hour road trip. If they all shared the drive time equally, how many hours did each family member drive? Answer two-thirds 1 and one-fifth five-sixths 1 and four-ninths

OpenStudy (anonymous):

\[3\frac{1}{3} = \frac{10}{3}\]\[\frac{10}{3} \div 4 = \frac{5}{6}\]

OpenStudy (anonymous):

Estimate 2 and six-sevenths times 5 and one-fourth by rounding each factor to the nearest whole number. Explain your choices for rounding. Please help.

OpenStudy (anonymous):

\[2\frac{6}{11} \times 5 \frac{1}{4}=\frac{28}{11} \times \frac{21}{4} = \frac{588}{44} \]\[= 13.36363636 \approx 13\]

OpenStudy (anonymous):

Multiply seven-eighths times twenty-eight

OpenStudy (anonymous):

\[\frac{7}{8} \times 28 = \frac{49}{2}\]